0 0
Add a comment Improve this question Transcribed image text
Answer #1

2 2 O.SY - 87,48 PC
Answer #3

Authenticated and Verified Answer from latest Electromagnetic literature:

Firstly , from Coulomb's Law we know that : F= kq1. q2/r^2

For the field developed by charge q=q1 at a distance r,

E = kq / r2

Now, simply substitute the valuesof k,r and E.

2.7 N/C = ( 9 x 109 x q ) / ( 0.542 )

q = 8.75 x 10-11 C

Hope this helps...please rate....thanx....
Answer #4
  • E = kq / r2
  • 2.7 N/C = ( 9 x 109 x q ) / ( 0.542 )
  • q = 8.75 x 10-11 C , UNIT = Coulumb